1 頭文字がでかい

This is a very simple sample file fordropping package. ‘Hoge’ is hoge ほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほ

げほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげほげ。 .